# Lusztig sheaves, characteristic cycles and the Borel-Moore homology of Nakajima’s quiver varieties

### v1

@@ 4.2.1. The functor $\mathcal{F}^{(r)}_{i}$
$\mathcal{F}^{(r)}_{i}$ $=\mathbf{Ind}^{\mathbf{V}\oplus\mathbf{W}^{1}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}}(\overline{\mathbb{Q}}_{l}|_{\mathbf{E}_{\mathbf{V}^{\prime},\Omega}}\boxtimes-):\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{\prime\prime}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1},\Omega^{(1)}})\rightarrow\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}),$
$\mathcal{F}^{(r)}_{i}$ $=\mathbf{Ind}^{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}(\overline{\mathbb{Q}}_{l}|_{\mathbf{E}_{\mathbf{V}^{\prime},\Omega}}\boxtimes-):\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{\prime\prime}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})\rightarrow\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}).$
@@ 4.2.2. The functor $\mathcal{E}^{(r)}_{i}$
Recall that the torus $\mathbb{T}$ of rank $\sharp\Omega^{(1)}$ acts on $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\mathbf{E}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1},\Omega^{(1)}}$ by scaling, and there are natural $\mathbb{T}$-actions on $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}^{i,0},\mathbf{E}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1},\Omega^{(1)}}^{i,0},\dot{\mathbf{E}}^{i}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\dot{\mathbf{E}}^{i}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1},\Omega^{(1)}},\mathbf{Grass}(\nu_{i},\tilde{\nu}_{i}),\mathbf{Flag}(\nu^{\prime\prime}_{i},\nu_{i},\tilde{\nu}_{i})$ such that all morphisms appearing in above diagram are $\mathbb{T}$-equivariant. We define the functor
@@
Above definitions of $\mathcal{E}^{(r)}_{i}$ use the assumption that $i$ is a source for the orientation $\Omega$. In general, for any orientation $\Omega$ and $i\in I$, we fix a orientation $\Omega_{i}$ such that $i$ is a source for $\Omega_{i}$, then we use $\mathcal{E}^{(r)}_{i}$ associated to $\Omega_{i}$ and Fourier-Laumon transforms to define the following functors
@@
Proposition 4.4 ([8, Proposition 3.18, Corollary 3.19]).
@@
The proof is essentially the same as the proof of [8, Proposition 3.18, Corollary 3.19]. ∎
@@ 4.2.3. The functors $\mathcal{K}_{i},\mathcal{K}_{-i}$
$\mathcal{K}_{i}={\mathrm{Id}}[-2\nu_{i}+\sum_{h\in\Omega,h^{\prime}=i}\nu_{h^{\prime\prime}}+\omega^{1}_{i}]:\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu}\rightarrow\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu},$
$\mathcal{K}_{-i}={\mathrm{Id}}[2\nu_{i}-\sum_{h\in\Omega,h^{\prime}=i}\nu_{h^{\prime\prime}}-\omega^{1}_{i}]:\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu}\rightarrow\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu},$
$\mathcal{K}_{i}={\mathrm{Id}}[-2\nu_{i}+\sum_{h\in\Omega,h^{\prime}=i}\nu_{h^{\prime\prime}}+\omega^{1}_{i}+\omega^{2}_{i}]:\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu}\rightarrow\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu},$
$\mathcal{K}_{-i}={\mathrm{Id}}[2\nu_{i}-\sum_{h\in\Omega,h^{\prime}=i}\nu_{h^{\prime\prime}}-\omega^{1}_{i}-\omega^{2}_{i}]:\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu}\rightarrow\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu}.$
@@ 4.3. Categorification theorem
send [¯$[\overline{\mathbb{Q}}_{l}|_{\mathbf{E}_{0\oplus\mathbf{W}^{1},\Omega^{(1)}}}]$ to the highest weight vector of ${{}_{\mathcal{A}}L}_{v}(\lambda_{1})$. Moreover, $\mathcal{L}(\omega^{1})$ has a $\mathcal{A}$-basis
@@
$\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})\xrightarrow{\mathbf{D}}\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})\xrightarrow{\mathbf{Res}^{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}_{\mathbf{V}^{2}\oplus 0\oplus\mathbf{W}^{2},\mathbf{V}^{1}\oplus\mathbf{W}^{1}\oplus 0}}$
$\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{2}}\times\mathbb{T}\times\mathbf{G}_{\mathbf{V}^{1}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{2}\oplus\mathbf{W}^{2},\Omega^{(1)}}\times\mathbf{E}_{\mathbf{V}^{1}\oplus\mathbf{W}^{1},\Omega^{(1)}})\xrightarrow{\mathbf{D}\boxtimes\mathbf{D}}\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{2}}\times\mathbb{T}\times\mathbf{G}_{\mathbf{V}^{1}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{2}\oplus\mathbf{W}^{2},\Omega^{(1)}}\times\mathbf{E}_{\mathbf{V}^{1}\oplus\mathbf{W}^{1},\Omega^{(1)}})\xrightarrow{\mathrm{sw}}$
$\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{1}}\times\mathbb{T}\times\mathbf{G}_{\mathbf{V}^{2}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{1}\oplus\mathbf{W}^{1},\Omega^{(1)}}\times\mathbf{E}_{\mathbf{V}^{2}\oplus\mathbf{W}^{2},\Omega^{(1)}}),$
where $\mathbf{D}$ is the Verdier dual functor, $\mathbf{Res}^{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}_{\mathbf{V}^{2}\oplus 0\oplus\mathbf{W}^{2},\mathbf{V}^{1}\oplus\mathbf{W}^{1}\oplus 0}$ is the Lusztig’s restriction functor for the two-framed quiver $Q^{(2)}$, and $\mathrm{sw}$ is the natural isomorphism induced by swapping two coordinates.
Theorem 4.6 ([7, Theorem 4.15]).
Above composition of functors induces an ${{}_{\mathcal{A}}\mathbf{U}_{v}}$-module isomorphism $\Delta:\mathcal{L}(\omega^{1},\omega^{2})\rightarrow\mathcal{L}(\omega^{1})\otimes\mathcal{L}(\omega^{2})$ such that $mL(\omega^{1},\omega^{2})$ is isomorphic to ${}_{\mathcal{A}}L_{v}(\lambda_{1})\otimes{{}_{\mathcal{A}}L}_{v}(\lambda_{2})$ via
$\tilde{\chi}^{\omega^{1},\omega^{2}}=(\chi^{\omega^{1}}\otimes\chi^{\omega^{2}})\Delta:mL(\omega^{1},\omega^{2})\rightarrow_{\mathcal{A}}L_{v}(\lambda_{1})\otimes{{}_{\mathcal{A}}L}_{v}(\lambda_{2}).$
@@
$\bigsqcup_{\nu\in\mathbb{N}I}\{\tilde{\chi}^{\omega^{1},\omega^{2}}([L])\in_{\mathcal{A}}L_{v}(\lambda_{1})\otimes{{}_{\mathcal{A}}L}_{v}(\lambda_{2})\mid L\in\mathcal{P}_{\nu,\omega^{1},\omega^{2}},[L]\not=0\ \textrm{in}\ \mathcal{L}(\nu,\omega^{1},\omega^{2})\}$
under the isomorphism $\tilde{\chi}^{\omega^{1},\omega^{2}}$ is the canonical basis $\mathcal{B}_{\lambda_{1},\lambda_{2}}$ of ${}_{\mathcal{A}}L_{v}(\lambda_{1})\otimes{{}_{\mathcal{A}}L}_{v}(\lambda_{2})$ defined in [2].
@@ 5.1.1. Lusztig’s nilpotent variety
such that $\mathbf{E}_{\mathbf{V},H}$ can be identified with the cotangent bundle of $\mathbf{E}_{\mathbf{V},\Omega}$, see [14, Section 12.8]. The moment map attached to the $\mathbf{G}_{\mathbf{V}}$-action on the symplectic vector space $\mathbf{E}_{\mathbf{V},H}$ is
@@
where $x\in\mathbf{E}_{\mathbf{V},H}$ is said to be nilpotent, if there exists an $N\geqslant 2$ such that for any sequence $h_{1},...,h_{N}\in H$ satisfying $h_{1}^{\prime}=h_{2}^{\prime\prime},...,h_{N-1}^{\prime}=h_{N}^{\prime\prime}$, the composition $x_{h_{1}}...x_{h_{N}}:\mathbf{V}_{h^{\prime}_{N}}\rightarrow\mathbf{V}_{h_{1}^{\prime\prime}}$ is zero. By [14, Theorem 12.9], the nilpotent variety $\Lambda_{\mathbf{V}}$ is a Lagrangian subvariety of $\mathbf{E}_{\mathbf{V},H}$.

### v3

@@ 4.2.1. The functor $\mathcal{F}^{(r)}_{i}$
$\mathcal{F}^{(r)}_{i}$ $=\mathbf{Ind}^{\mathbf{V}\oplus\mathbf{W}^{1}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}}(\mathbb{C}|_{\mathbf{E}_{\mathbf{V}^{\prime},\Omega}}\boxtimes-):\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{\prime\prime}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1},\Omega^{(1)}})\rightarrow\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}),$
$\mathcal{F}^{(r)}_{i}$ $=\mathbf{Ind}^{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}(\mathbb{C}|_{\mathbf{E}_{\mathbf{V}^{\prime},\Omega}}\boxtimes-):\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{\prime\prime}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})\rightarrow\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}).$
@@ 4.2.2. The functor $\mathcal{E}^{(r)}_{i}$
Recall that the torus $\mathbb{T}$ of rank $\sharp I$ acts on $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\mathbf{E}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1},\Omega^{(1)}}$ by scaling, and there are natural $\mathbb{T}$-actions on $\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}^{i,0},\mathbf{E}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1},\Omega^{(1)}}^{i,0},\dot{\mathbf{E}}^{i}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\dot{\mathbf{E}}^{i}_{\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1},\Omega^{(1)}},\mathbf{Grass}(\nu_{i},\tilde{\nu}_{i}),\mathbf{Flag}(\nu^{\prime\prime}_{i},\nu_{i},\tilde{\nu}_{i})$ such that all morphisms appearing in above diagram are $\mathbb{T}$-equivariant. We define the functor
@@
Above definitions of $\mathcal{E}^{(r)}_{i}$ use the assumption that $i$ is a source for the orientation $\Omega$. In general, for any orientation $\Omega$ and $i\in I$, we fix a orientation $\Omega_{i}$ such that $i$ is a source for $\Omega_{i}$ and $\Omega^{(N)},\Omega^{(N)}_{i}$ then we use $\mathcal{E}^{(r)}_{i}$ associated to $\Omega_{i}$ and Fourier-Sato transforms to define the following functors
@@
Proposition 4.4 ([12, Proposition 3.18, Corollary 3.19]).
@@
The proof is essentially the same as the proof of [12, Proposition 3.18, Corollary 3.19]. ∎
@@ 4.2.3. The functors $\mathcal{K}_{i},\mathcal{K}_{-i}$
$\mathcal{K}_{i}={\mathrm{Id}}[-2\nu_{i}+\sum_{h\in\Omega_{i},h^{\prime}=i}\nu_{h^{\prime\prime}}+\omega^{1}_{i}]:\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu}\rightarrow\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu},$
$\mathcal{K}_{-i}={\mathrm{Id}}[2\nu_{i}-\sum_{h\in\Omega_{i},h^{\prime}=i}\nu_{h^{\prime\prime}}-\omega^{1}_{i}]:\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu}\rightarrow\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu},$
$\mathcal{K}_{i}={\mathrm{Id}}[-2\nu_{i}+\sum_{h\in\Omega_{i},h^{\prime}=i}\nu_{h^{\prime\prime}}+\omega^{1}_{i}+\omega^{2}_{i}]:\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu}\rightarrow\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu},$
$\mathcal{K}_{-i}={\mathrm{Id}}[2\nu_{i}-\sum_{h\in\Omega_{i},h^{\prime}=i}\nu_{h^{\prime\prime}}-\omega^{1}_{i}-\omega^{2}_{i}]:\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu}\rightarrow\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu}.$
@@ 4.3. Categorification theorem
send $[\mathbb{C}|_{\mathbf{E}_{0\oplus\mathbf{W}^{1},\Omega^{(1)}}}]$ to the highest weight vector of ${{}_{\mathcal{A}}L}_{v}(\lambda_{1})$. Moreover, $\mathcal{L}(\omega^{1})$ has a $\mathcal{A}$-basis
@@
$\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})\xrightarrow{\mathbf{D}}\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})\xrightarrow{\mathbf{Res}^{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}_{\mathbf{V}^{1}\oplus\mathbf{W}^{1},\mathbf{V}^{2}\oplus\mathbf{W}^{2}}}$
$\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{1}}\times\mathbb{T}\times\mathbf{G}_{\mathbf{V}^{2}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{1}\oplus\mathbf{W}^{1},\Omega^{(1)}}\times\mathbf{E}_{\mathbf{V}^{2}\oplus\mathbf{W}^{2},\Omega^{(1)}})\xrightarrow{\mathbf{D}\boxtimes\mathbf{D}}\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{1}}\times\mathbb{T}\times\mathbf{G}_{\mathbf{V}^{2}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{1}\oplus\mathbf{W}^{1},\Omega^{(1)}}\times\mathbf{E}_{\mathbf{V}^{2}\oplus\mathbf{W}^{2},\Omega^{(1)}})\xrightarrow{\mathrm{sw}}$
$\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{2}}\times\mathbb{T}\times\mathbf{G}_{\mathbf{V}^{1}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{2}\oplus\mathbf{W}^{2},\Omega^{(1)}}\times\mathbf{E}_{\mathbf{V}^{1}\oplus\mathbf{W}^{1},\Omega^{(1)}}),$
where $\mathbf{D}$ is the Verdier dual functor, $\mathbf{Res}^{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}}_{\mathbf{V}^{1}\oplus\mathbf{W}^{1},\mathbf{V}^{2}\oplus\mathbf{W}^{2}}$ is the Lusztig’s restriction functor for the two-framed quiver $Q^{(2)}$, and $\mathrm{sw}$ is the natural isomorphism induced by swapping two coordinates.
Theorem 4.6 ([10, Theorem 4.15]).
Above composition of functors induces an ${{}_{\mathcal{A}}\mathbf{U}_{v}}$-module isomorphism $\Delta:\mathcal{L}(\omega^{1},\omega^{2})\rightarrow\mathcal{L}(\omega^{1})\otimes\mathcal{L}(\omega^{2})$ such that $\mathcal{L}(\omega^{1},\omega^{2})$ is isomorphic to ${}_{\mathcal{A}}L_{v}(\lambda_{2})\otimes{{}_{\mathcal{A}}L}_{v}(\lambda_{1})$ via
$\tilde{\chi}^{\omega^{1},\omega^{2}}=(\chi^{\omega^{1}}\otimes\chi^{\omega^{2}})\Delta:\mathcal{L}(\omega^{1},\omega^{2})\rightarrow_{\mathcal{A}}L_{v}(\lambda_{2})\otimes{{}_{\mathcal{A}}L}_{v}(\lambda_{1}).$
@@
$\bigsqcup_{\nu\in\mathbb{N}I}\{\tilde{\chi}^{\omega^{1},\omega^{2}}([L])\in_{\mathcal{A}}L_{v}(\lambda_{2})\otimes{{}_{\mathcal{A}}L}_{v}(\lambda_{1})\mid L\in\mathcal{P}_{\nu,\omega^{1},\omega^{2}},[L]\not=0\ \textrm{in}\ \mathcal{L}(\nu,\omega^{1},\omega^{2})\}$
under the isomorphism $\tilde{\chi}^{\omega^{1},\omega^{2}}$ is the canonical basis $\mathcal{B}_{\lambda_{1},\lambda_{2}}$ of ${}_{\mathcal{A}}L_{v}(\lambda_{2})\otimes{{}_{\mathcal{A}}L}_{v}(\lambda_{1})$ defined in [3].
Remark 4.7.
In Lemma 3.3, we only consider the Fourier-Sato transforms for derived categories associated to equivalent orientations. More generally, we can consider the vector bundle
$\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}\rightarrow\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}\cap\Omega^{(1),^{\prime}}},$
$\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}\rightarrow\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}\cap\Omega^{(2),^{\prime}}}$
for any orientations $\Omega^{(N)},\Omega^{(N),^{\prime}}$. All Lusztig’s semisimple sheaves are monodromic with respect to this vector bundle, and the Lusztig’s sheaves in thick subcategories $\mathcal{N}_{\nu}$ are preserved under the Fourier-Sato transform. It implies that our construction is independent on the choice of the orientation, but not only independent on the choice of mutation equivalent class of the orientation.
@@ 5.1.1. Lusztig’s nilpotent variety
such that $\mathbf{E}_{\mathbf{V},H}$ can be identified with the cotangent bundle of $\mathbf{E}_{\mathbf{V},\Omega}$, see [18, Section 12.8]. The moment map attached to the $\mathbf{G}_{\mathbf{V}}$-action on the symplectic vector space $\mathbf{E}_{\mathbf{V},H}$ is
@@
where $x\in\mathbf{E}_{\mathbf{V},H}$ is said to be nilpotent, if there exists an $N\geqslant 2$ such that for any sequence $h_{1},...,h_{N}\in H$ satisfying $h_{1}^{\prime}=h_{2}^{\prime\prime},...,h_{N-1}^{\prime}=h_{N}^{\prime\prime}$, the composition $x_{h_{1}}...x_{h_{N}}:\mathbf{V}_{h^{\prime}_{N}}\rightarrow\mathbf{V}_{h_{1}^{\prime\prime}}$ is zero. By [18, Theorem 12.9], the nilpotent variety $\Lambda_{\mathbf{V}}$ is a Lagrangian subvariety of $\mathbf{E}_{\mathbf{V},H}$.